You are the manager of a firm that produces furniture. All your clients are in Kumasi (K) and Accra (A). Suppose the monthly inverse demand function for your furniture in in Accra and Kumasi combined is 𝑃 = 1200 βˆ’ 4𝑄. Because of prohibitive cost of transporting furniture between Accra and Kumasi, you set up two plants, one in each town. The cost of producing furniture at each facility is given by 𝐢A (𝑄A ) = 8000 + 6𝑄A2 and 𝐢K(𝑄K) = 8000 + 6𝑄K2 where 𝑄 = 𝑄A + 𝑄K and 𝑄A and 𝑄K are the quantity of furniture produced at Accra and Kumasi respectively. Determine the profit-maximizing amounts of electricity to produce in the two facilities, the optimal price and the profit.


1
Expert's answer
2021-09-28T13:17:23-0400

Solution:

Profit maximizing level of output: MR = MC

Derive MR:

TR = P x Q = (1,200 – 4Q) Q = 1,200Q – 4Q2

TR = 1,200Q – 4Q2

MR = "\frac{\partial TR} {\partial Q}" = 1,200 – 8Q

MR = 1,200 – 8QA – 8QK

Β 

MR = MC1 and MR = MC2

TC1 (QA) = 8000 + 6QA2

MC1 (QA) = "\frac{\partial TC_{1} } {\partial Q_{A} }" = 12QA


TC2 (QK) = 8000 + 6QK2

MC2 (QK) "\frac{\partial TC_{2} } {\partial Q_{K} }" = 12QK

Q = QA + QK

Set the two equations:

MC1 (QA) = MR

MC2 (QK) = MR

Β 

1). 1,200 – 8QA – 8QK = 12QA

1200 – 8QK = 12QA + 8QA

1200 – 8QK = 20QA

Divide all sides by 20:

60 – 0.4QK = QA

QA = 60 – 0.4QK

Substitute for QK in the second function:

2.). 1200 – 8QA – 8QK = 12QK

1200 – 8(60 – 0.4QK) – 8QK = 12QK

1200 – 480 + 3.2QK – 8QK = 12QK

720 = 12QK + 8QK – 3.2QK

720 = 16.8QK

QK = 43

QA = 60 – 0.4QK = 60 – 0.4(43) = 60 – 17 = 43

QA = 43

Q = 43 + 43 = 86

The profit-maximizing amounts of electricity to produce in the two facilities = 86 Units


The optimal price:

P = 1200 – 4Q = 1200 – 4(86) = 1200 – 344 = 856

The optimal price = 856

Β 

Profit = TR – TC

TR = P "\times" Q = 856 "\times" 86 = 73,616

TC = TC1 + TC2

TC1 = 8000 + 6𝑄A2Β = 8000 + 6(432) = 8000 + 11094 = 19,094

TC2 = = 8000 + 6𝑄K2Β = 8000 + 6(432) = 8000 + 11094 = 19,094

TC = 19,094 + 19,094 = 38,188


Profit = 73,616 – 38,188 = 35,428

Profit = 35,428